# Movable Antenna Empowered Covert Dual-Functional Radar-Communication

### v1

@@ I Introduction
The burgeoning demand for intelligent and ubiquitous services, such as smart grids, unmanned aerial vehicles (UAVs), and the Industrial Internet of Things (IIoT), has become a primary catalyst for the evolution toward sixth-generation (6G) mobile networks. These emerging applications necessitate unprecedented performance metrics, including a peak data rate of 1 Tb/s, a user-experienced data rate of 10–100 Gb/s, and centimetric localization accuracy [2]. To fulfill these stringent requirements, dual-functional radar-communication (DFRC) has emerged as a pivotal transformative technology within the International Mobile Telecommunication (IMT)-2030 framework [3]. By integrating radar sensing and communication functions in a unified platform, DFRC enables dual functions to share spectrum resources, hardware facilities, and signal-processing modules, leading to significantly improved system capacity and resource utilization efficiency [4]. However, the broadcast nature of wireless channels poses significant security vulnerabilities. Specifically, in DFRC systems, allowing unified probing waveforms to carry private information will cause a high security risk of being wiretapped if the sensing targets are malicious eavesdroppers [5, 6]. Consequently, the development of efficient security solutions for DFRC systems is highly imperative.
@@
To fulfill the demands for ever-increasing security requirements, covert communication, which shields confidential transmission behaviors from Willies, has been proposed to provide a higher level of security[13]. The achievability of the square root law was first established in [14] where Alice can covertly transmit a maximum of $\mathcal{O}(\sqrt{n})$ bits of information to Bob over $n$ channel uses. The authors in [15] further proposed that $\mathcal{O}(n)$ bits can be covertly transmitted with the aid of AN. To overcome these information-theoretic limitations, various advanced technologies are employed to increase the covert transmission rate in DFRC systems [16, 17, 18, 19, 20]. For instance, in [16], a multi-strategy alternating optimization framework was proposed to enhance transmission covertness for reconfigurable intelligent surface (RIS)-enhanced DFRC systems. The authors in [17] studied the covert DFRC systems against multiple randomly distributed Willies. A robust beamforming optimization model was studied in [19] to ensure data covertness. Although these works have demonstrated the effectiveness of covert communications, the inherent existence of wireless fading caused by complex propagation environment severely undermines covertness performance. Furthermore, existing literature mainly focused on the transceiver design by employing conventional fixed-position antennas (FPAs), while channel variations in the continuous spatial field were not fully exploited. Additionally, the fixed geometric configurations of FPA arrays can result in array-gain loss during dual-functional beamforming, leading to degraded communication security.
@@
To overcome the bottleneck of conventional FPA-based systems, movable antennas (MAs), or fluid antennas, have recently been proposed as a promising solution to enhance the dual-task performance [21, 22]. In MA-assisted systems, each antenna element is connected to a radio frequency (RF) chain via flexible cables to support active antenna movement. A prototype of the MA-assisted radar system was initially demonstrated in [23], while the latest version achieved an antenna positioning accuracy of up to 0.05 mm in [24]. Channel modeling and performance analysis were explored under far-field channel conditions in [25]. Based on the results in [25], a few works have investigated secure designs for MA-enhanced DFRC systems [26, 27, 28, 29]. However, the majority of existing literature, i.e., [26, 27, 28] and the references therein, primarily focused on safeguarding confidential information from being deciphered instead of concealing the transmission itself, which deteriorates data covertness. To achieve higher security, the authors in [29] pioneered the investigation of covert transmission in MA-enabled DFRC systems. Unfortunately, the proposed approach is limited to scenarios with a single radar target and a solitary Willie. To the best of our knowledge, the potential of MAs for enhancing transmission covertness within DFRC systems remains largely under-explored, particularly in more complex and practical scenarios. To further bolster security performance, a unified covert design framework is highly desirable.
@@ II System Model
We consider a narrowband DFRC system, as depicted in Fig. 1, where a dual-functional base station (BS) serves $K$ covert users (Bobs) while simultaneously sensing $W$ point-like targets. The BS is equipped with two separate MA-based uniform linear arrays (ULAs), each consisting of $N$ movable antennas, dedicated to signal transmission and reception, respectively. The sensing targets are assumed to be malicious Willies that attempt to detect whether the BS is transmitting confidential information or not. We exploit dedicated radar signals as a cover to achieve covert communications. We assume that the feasible movement range for both the transmitting and receiving MAs is a one-dimensional (1D) interval of length $D$. The transmitting and receiving antenna positioning vectors (APVs) are denoted by $\bm{t}=[t_{1},t_{2},\dots,t_{N}]^{T}\in\mathbb{R}^{N\times 1}$ and $\bm{r}=[r_{1},r_{2},\dots,r_{N}]^{T}\in\mathbb{R}^{N\times 1}$, respectively, with $0\leq t_{1}\leq t_{2}\dots\leq t_{N}\leq D$ and $0\leq r_{1}\leq r_{2}\dots\leq r_{N}\leq D$.
@@ II-A Signal Model
Denote by $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ the hypotheses that the BS transmits covert signals or not, respectively. The transmitted signal can be given by
@@
Here, $\bm{s}(m)=[s_{1}(m),s_{2}(m),\dots,s_{K}(m)]^{T}\in\mathbb{C}^{K\times 1}$ denotes the communication symbols for $K$ covert users in the $m$-th time slot, $\forall m\in\mathcal{M}=\{1,\dots,M\}$. Meanwhile, $\bm{W}=[\bm{w}_{1},\dots,\bm{w}_{K}]\in\mathbb{C}^{N\times K}$ denotes the transmitting beamforming matrix, and $\bm{r}(m)\in\mathbb{C}^{N\times 1}$ is the dedicated radar signal. Following the work in [30, 31], we assume that both $\bm{s}(m)$ and $\bm{r}(m)$ are zero-mean, temporally-white, and wide-sense stationary stochastic processes, and the symbols are uncorrelated between different users. It is assumed that $\bm{s}(m)$ and $\bm{r}(m)$ are independent Gaussian distributed with $\bm{s}(m)\sim\mathcal{CN}(\bm{0},\bm{I}_{K})$ and $\bm{r}(m)\sim\mathcal{CN}(\bm{0},\bm{R}_{0})$, respectively, where $\bm{R}_{0}\in\mathbb{C}^{N\times N}$ is the covariance matrix of a general rank due to potential multiple beam transmission. We note that once $\bm{R}_{0}$ is determined, the dedicated radar signal $\bm{r}(m)$ can be generated [32]. Consequently, the covariance matrix of the transmitted signal $\bm{x}(m)$ can be derived as
$\begin{cases}\mathcal{H}_{0}:\bm{R}_{X}^{0}=\bm{R}_{0},\;\;\\ \mathcal{H}_{1}:\bm{R}_{X}^{1}=\sum_{k=1}^{K}\bm{w}_{k}\bm{w}_{k}^{H}+\bm{R}_{0}.\end{cases}$ (2)
@@ II-C Radar Model
III Problem Formulation
In this section, we first formulate a general covert rate maximization problem for MA-enhanced DFRC systems. Subsequently, we elaborate on the optimal detection performance of Willies under different detection modes, and derive tractable covertness constraints.
III-A Problem Formulation
@@
$\hphantom{s.t.~}\xi^{\star}\geq 1-\epsilon,$ (10g)
where $\Gamma_{w}$ is the radar SINR threshold for the $w$-th target, $d$ represents the minimum distance between MAs to prevent coupling effects, and $P_{t}$ is the maximum transmission power. Note that the constraint in (10g) characterizes a general covertness requirement, where $\xi^{\star}$ denotes the minimum detection error probability at Willies, and $\epsilon\in[0,1]$ specifies the covertness level of confidential transmission.
The problem in (10) is computationally intractable due to the highly non-convex objective function and tight coupling among optimization variables. The implicit dependence of Willies’ DEP on beamforming vectors and antenna placement across different detection modes further exacerbates the difficulty of algorithm development. To facilitate system design, we first derive the optimal detectors and corresponding minimum DEP for Willies in both non-colluding and colluding scenarios. Based on these analytical results, we then develop tractable covertness constraints to simplify the optimization.
III-B Covertness Constraints for Non-colluding Willies
Recall that the sensing targets are assumed to be malicious Willies who aim to distinguish between the two hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ based on received signals. Under non-colluding scenarios, each Willie performs detection independently. Specifically, the signals received at Willie $w$ over $M$ time slots in one round are stacked into an observation vector, given by $\bm{y}_{w}=[y_{w}(1),y_{w}(2),\dots,y_{w}(M)]^{T}\in\mathbb{C}^{M\times 1}$, where each element $y_{w}(m)$ is expressed as
Theorem 1: The optimal test for Willie $w$ to detect the covert communication behaviours is given by
$\bm{y}_{w}||^{2}_{2}\overset{\mathcal{D}_{1}}{\underset{\mathcal{D}_{0}}{\gtrless}}M\frac{\eta_{w,0}\eta_{w,1}}{\eta_{w,1}-\eta_{w,0}}\ln\frac{\eta_{w,1}}{\eta_{w,0}}\triangleq\varpi^{*}_{w},$ (12)
where $\bm{y}_{w}||^{2}_{2}$ denotes the received signal power at Willie $w$, $\eta_{w,0}=|\beta_{w}|^{2}\mathbf{a}_{t}(\varphi_{w},{\bm{t}})^{H}\bm{R}_{0}\mathbf{a}_{t}(\varphi_{w},{\bm{t}})+\sigma_{0}^{2}$, and $\eta_{w,1}=\sum_{k=1}^{K}|\beta_{w}|^{2}|\mathbf{a}_{t}(\varphi_{w},{\bm{t}})^{H}\bm{w}_{k}|^{2}+\eta_{w,0}$. Consequently, the minimum DEP $\xi^{\star}_{w}$ is given by
@@ III-B Covertness Constraints for Non-colluding Willies
[…]
Proof: Please refer to [1]. $\Box$
With a given covertness level $\epsilon_{w}$ against Willie $w$, directly addressing the covertness constraint $\xi^{\star}_{w}\geq 1-\epsilon_{w}$ is highly intractable due to the lower incomplete Gamma functions. Fortunately, according to Pinsker’s inequality in [13], we have $\mathcal{V}_{T}(\mathbb{P}_{w,0},\mathbb{P}_{w,1})\leq\sqrt{\frac{1}{2}\mathcal{D}(\mathbb{P}_{w,0}||\mathbb{P}_{w,1})}$, where $\mathcal{D}(\mathbb{P}_{w,0}||\mathbb{P}_{w,1})$ is the Kullback-Leibler (KL) divergence from $\mathbb{P}_{w,0}$ to $\mathbb{P}_{w,1}$, given by
Therefore, the generalized constraint in (10g) can be specified as
$\mathcal{D}(\mathbb{P}_{w,0}||\mathbb{P}_{w,1})\leq 2\epsilon^{2}_{w},\forall w\in\mathcal{W},$ (15)
where a smaller $\epsilon_{w}$ corresponds to a more stringent covertness requirement for Willie $w$.
III-C Covertness Constraints for Colluding Willies
$\bm{y}_{F}(m)=\bm{H}_{F}(\bm{t})\bm{x}(m)+\bm{n}_{F}(m),$ (16)
where $\lambda_{1}^{0},\dots,\lambda_{W}^{0}$ and ${\lambda}_{1}^{1},\dots,{\lambda}_{W}^{1}$ are the eigenvalues of $\bm{\Lambda}_{0}^{1/2}\bm{V}\bm{\Lambda}_{0}^{1/2}$ and $\bm{\Lambda}_{1}^{1/2}\bm{V}\bm{\Lambda}_{1}^{1/2}$, respectively. $F(\chi|\frac{1}{{\lambda}_{1}^{i}},\dots,\frac{1}{{\lambda}_{W}^{i}})$ is defined in (19), shown at the top of this page. Besides, $\mathcal{V}_{T}(\mathbb{P}_{F,0},\mathbb{P}_{F,1})$ is the total variation between PDFs of $\bm{Y}_{F}$ under $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, where $\mathbb{P}_{F,i}$ is elaborated in (51).
Proof: Please refer to Appendix A. $\Box$
The expression of the theoretical minimum DEP in (18) is rather intractable to guide further system design. According to Pinsker’s inequality in [13], $\mathcal{V}_{T}(\mathbb{P}_{F,0},\mathbb{P}_{F,1})$ is upper bounded by $\mathcal{V}_{T}(\mathbb{P}_{F,0},\mathbb{P}_{F,1})\leq\sqrt{\frac{1}{2}\mathcal{D}(\mathbb{P}_{F,0}||\mathbb{P}_{F,1})}$, where the KL divergence $\mathcal{D}(\mathbb{P}_{F,0}||\mathbb{P}_{F,1})$ is derived as
As such, with a given covertness level $\epsilon_{F}$ against colluding Willies, the generalized constraint in (10g) can be specified as
$\mathcal{D}(\mathbb{P}_{F,0}||\mathbb{P}_{F,1})\leq 2\epsilon_{F}^{2}.$ (21)
IV Proposed BCD Algorithm

### v2

@@ I Introduction
The burgeoning demand for intelligent and ubiquitous services, such as smart grids, unmanned aerial vehicles (UAVs), and the Industrial Internet of Things (IIoT), has become a primary catalyst for the evolution toward sixth-generation (6G) mobile networks. These emerging applications necessitate unprecedented performance metrics, including a peak data rate of 1 Tb/s, a user-experienced data rate of 10–100 Gb/s, and centimetric localization accuracy [2]. To fulfill these stringent requirements, dual-functional radar-communication (DFRC) has emerged as a pivotal transformative technology within the International Mobile Telecommunication (IMT)-2030 framework [3]. In particular, DFRC enables dual functions to share spectrum resources, hardware facilities, and signal-processing modules, leading to significantly improved system capacity and resource utilization efficiency [4], thanks to the integration of radar sensing and communication functions in a unified platform. However, the broadcast nature of wireless channels poses significant security vulnerabilities. Specifically, in DFRC systems, allowing unified probing waveforms to carry private information will cause a high security risk of being wiretapped if the sensing targets are malicious eavesdroppers [5, 6]. Consequently, the development of efficient security solutions for DFRC systems is highly imperative.
@@
To fulfill the demands for ever-increasing security requirements, covert communication, which shields confidential transmission behaviors from Willies, has been proposed to provide a higher level of security[13]. The achievability of the square root law was first established in [14] where Alice can covertly transmit a maximum of $\mathcal{O}(\sqrt{n})$ bits of information to Bob over $n$ channel uses. The authors in [15] further proposed that $\mathcal{O}(n)$ bits can be covertly transmitted with the aid of AN. To overcome these information-theoretic limitations, various advanced technologies are employed to increase the covert transmission rate in DFRC systems [16, 17, 18, 19, 20]. For instance, in [16], a multi-strategy alternating optimization framework was proposed to enhance transmission covertness for reconfigurable intelligent surface (RIS)-enhanced DFRC systems. The authors in [17] studied covert DFRC systems against multiple randomly distributed Willies. A robust beamforming optimization model was studied in [19] to ensure data covertness. Although these works have demonstrated the effectiveness of covert communications, the inherent existence of wireless fading caused by complex propagation environment severely undermines covertness performance. Furthermore, existing literature mainly focused on the transceiver design by employing conventional fixed-position antennas (FPAs), while channel variations in the continuous spatial field were not fully exploited. Additionally, the rigid array manifold of FPAs fails to mitigate the spatial conflict between sensing and communication directions, leading to unavoidable array-gain loss and degraded security.
@@
To overcome the bottleneck of conventional FPA-based systems, movable antennas (MAs), or fluid antennas, have recently been proposed as a promising solution to enhance the dual-task performance [21, 22, 23, 24, 25, 26]. In MA-assisted systems, each antenna element is connected to a radio frequency (RF) chain via flexible cables to support active antenna movement. A prototype of the MA-assisted radar system was initially demonstrated in [27], while the latest version achieved an antenna positioning accuracy of up to 0.05 mm in [28]. Channel modeling and performance analysis were explored under far-field channel conditions in [29]. Based on the results in [29], a few works have investigated secure designs for MA-enhanced DFRC systems [30, 31, 32, 33]. However, the majority of existing literature, i.e., [30, 31, 32] and the references therein, primarily focused on safeguarding confidential information from being deciphered instead of concealing the transmission itself, which deteriorates data covertness. To achieve higher security, the authors in [33] pioneered the investigation of covert transmission in MA-enabled DFRC systems. Unfortunately, the proposed approach is limited to scenarios with a single radar target and a solitary Willie. To the best of our knowledge, the potential of MAs for enhancing transmission covertness within DFRC systems remains largely under-explored, particularly in more complex and practical scenarios. To further bolster security performance, a unified covert design framework is highly desirable.
@@ II System Model
We consider a narrowband DFRC system, as depicted in Fig. 1, where a dual-functional base station (BS) serves $K$ covert users (Bobs) while simultaneously sensing $W$ point-like targets. The BS is equipped with two separate MA-based uniform linear arrays (ULAs), each consisting of $N$ movable antennas, dedicated to signal transmission and reception, respectively [^1] We assume that self-interference (SI) between transmit and receive antenna arrays is effectively suppressed by advanced SI cancellation methods, such as physical isolation and digital cancellation [22].. The sensing targets are assumed to be malicious Willies that attempt to detect whether the BS is transmitting confidential information or not. We exploit dedicated radar signals as a cover to achieve covert communications. We assume that the feasible movement range for both the transmitting and receiving MAs is a one-dimensional (1D) interval of length $D$. The transmitting and receiving antenna positioning vectors (APVs) are denoted by $\bm{t}=[t_{1},t_{2},\dots,t_{N}]^{T}\in\mathbb{R}^{N\times 1}$ and $\bm{r}=[r_{1},r_{2},\dots,r_{N}]^{T}\in\mathbb{R}^{N\times 1}$, respectively, with $0\leq t_{1}\leq t_{2}\dots\leq t_{N}\leq D$ and $0\leq r_{1}\leq r_{2}\dots\leq r_{N}\leq D$.
@@ II-A Signal Model
Denote by $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ hypotheses that the BS doesn’t transmit covert signals & does transmit covert signals, respectively. The transmitted signal can be given by
@@
Here, $\bm{s}(m)=[s_{1}(m),s_{2}(m),\dots,s_{K}(m)]^{T}\in\mathbb{C}^{K\times 1}$ denotes the communication symbols for $K$ covert users in the $m$-th time slot, $\forall m\in\mathcal{M}=\{1,\dots,M\}$. Meanwhile, $\bm{W}=[\bm{w}_{1},\dots,\bm{w}_{K}]\in\mathbb{C}^{N\times K}$ denotes the transmitting beamforming matrix, and $\bm{r}(m)\in\mathbb{C}^{N\times 1}$ is the dedicated radar signal. Following the work in [34, 35], we assume that both $\bm{s}(m)$ and $\bm{r}(m)$ are zero-mean, temporally-white, and wide-sense stationary stochastic processes, and the symbols are uncorrelated between different users. It is assumed that $\bm{s}(m)$ and $\bm{r}(m)$ are independently Gaussian distributed with $\bm{s}(m)\sim\mathcal{CN}(\bm{0},\bm{I}_{K})$ and $\bm{r}(m)\sim\mathcal{CN}(\bm{0},\bm{R}_{0})$, respectively, where $\bm{R}_{0}\in\mathbb{C}^{N\times N}$ is the covariance matrix of a general rank due to potential multiple beam transmission. We note that once $\bm{R}_{0}$ is determined, the dedicated radar signal $\bm{r}(m)$ can be generated [36]. Consequently, the covariance matrix of the transmitted signal $\bm{x}(m)$ can be derived as
$\begin{cases}\mathcal{H}_{0}:\bm{R}_{X}^{0}=\bm{R}_{0},\;\;\\ \mathcal{H}_{1}:\bm{R}_{X}^{1}=\sum_{k=1}^{K}\bm{w}_{k}\bm{w}_{k}^{H}+\bm{R}_{0},\end{cases}$ (2)
where $\bm{R}_{X}^{i},\forall i\in\{0,1\}$ denotes the covariance matrix of transmitted signals $\bm{x}(m)$ under $\mathcal{H}_{i}$.
@@ II-C Radar Model
II-D Detection Performance and Covertness Constraints
Recall that the sensing targets are assumed to be malicious Willies who aim to distinguish between the two hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ based on received signals. In the following, we develop the optimal detection criteria and covertness constraints for both non-colluding and colluding Willies, respectively.
II-D1 Non-colluding Willies
Under non-colluding scenarios, each Willie performs detection independently. Specifically, the signals received at Willie $w$ over $M$ time slots in one round are stacked into an observation vector, given by $\bm{y}_{w}=[y_{w}(1),y_{w}(2),\dots,y_{w}(M)]^{T}\in\mathbb{C}^{M\times 1}$, where each element $y_{w}(m)$ is expressed as
Theorem 1: The optimal test for Willie $w$ to detect the covert communication behaviours is derived as
[…]
where $\gamma(\cdot,\cdot)$ is the lower incomplete Gamma function given by $\gamma(M,x)=\int_{0}^{x}e^{-t}t^{M-1}dt$, and $\Gamma(M)=(M-1)!$ is the Gamma function. $\mathcal{V}_{T}(\mathbb{P}_{w,0},\mathbb{P}_{w,1})$ is the total variation between PDFs $\mathbb{P}_{w,0}$ and $\mathbb{P}_{w,1}$.
Based on the detection performance in (12), the covertness constraint for the $w$-th Willie can be given by
$\xi^{\star}_{w}\geq 1-\epsilon_{w},\forall w\in\mathcal{W},$ (13)
where $\epsilon_{w}$ is the given covertness level against Willie $w$ [14]. Note that a smaller $\epsilon_{w}$ corresponds to a more stringent covertness requirement for Willie $w$. Directly addressing the covertness constraint in (13) is highly intractable due to the lower incomplete Gamma functions. Fortunately, according to Pinsker’s inequality in [13], we have $\mathcal{V}_{T}(\mathbb{P}_{w,0},\mathbb{P}_{w,1})\leq\sqrt{\frac{1}{2}\mathcal{D}(\mathbb{P}_{w,0}||\mathbb{P}_{w,1})}$, where $\mathcal{D}(\mathbb{P}_{w,0}||\mathbb{P}_{w,1})$ is the Kullback-Leibler (KL) divergence from $\mathbb{P}_{w,0}$ to $\mathbb{P}_{w,1}$, given by
Therefore, the constraints in (13) can be approximated as
$\mathcal{D}(\mathbb{P}_{w,0}||\mathbb{P}_{w,1})\leq 2\epsilon^{2}_{w},\forall w\in\mathcal{W}.$ (15)
II-D2 Colluding Willies
$\bm{y}_{F}(m)=\bm{H}_{F}(\bm{t})\bm{x}(m)+\bm{n}_{F}(m),$ (16)
where $\lambda_{1}^{0},\dots,\lambda_{W}^{0}$ and ${\lambda}_{1}^{1},\dots,{\lambda}_{W}^{1}$ are the eigenvalues of $\bm{\Lambda}_{0}^{1/2}\bm{V}\bm{\Lambda}_{0}^{1/2}$ and $\bm{\Lambda}_{1}^{1/2}\bm{V}\bm{\Lambda}_{1}^{1/2}$, respectively. $F(\chi|\frac{1}{{\lambda}_{1}^{i}},\dots,\frac{1}{{\lambda}_{W}^{i}})$ is defined in (19), shown at the top of next page. Besides, $\mathcal{V}_{T}(\mathbb{P}_{F,0},\mathbb{P}_{F,1})$ is the total variation between PDFs of $\bm{Y}_{F}$ under $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, where $\mathbb{P}_{F,i},\forall i\in\{0,1\}$ is elaborated in (53).
Proof: Please refer to Appendix A. $\Box$
Similar to (13), the covertness constraint for colluding Willies can be given by
${\xi}_{F}^{\star}\geq 1-\epsilon_{F},$ (20)
where $\epsilon_{F}$ is the covertness level against colluding detection [14]. The expression of the theoretical minimum DEP in (18) is rather intractable to guide further system design. Similar to the non-colluding case, $\mathcal{V}_{T}(\mathbb{P}_{F,0},\mathbb{P}_{F,1})$ is upper bounded by $\mathcal{V}_{T}(\mathbb{P}_{F,0},\mathbb{P}_{F,1})\leq\sqrt{\frac{1}{2}\mathcal{D}(\mathbb{P}_{F,0}||\mathbb{P}_{F,1})}$, where the KL divergence $\mathcal{D}(\mathbb{P}_{F,0}||\mathbb{P}_{F,1})$ is derived as
As such, the constraint in (20) can be approximated as
$\mathcal{D}(\mathbb{P}_{F,0}||\mathbb{P}_{F,1})\leq 2\epsilon_{F}^{2}.$ (22)
II-E Problem Formulation
@@
where $\Gamma_{w}$ is the radar SINR threshold for the $w$-th target, $d$ represents the minimum distance between MAs to prevent coupling effects, $P_{t}$ is the maximum transmission power, and the constraint in (23g) guarantee the covertness level of confidential transmission under non-colluding or colluding Willie scenarios. We note that the problem in (23) is computationally intractable due to the highly non-convex objective function and tight coupling among optimization variables. The dependence of KL divergence on beamforming vectors and antenna placement across different detection modes further exacerbates the difficulty of algorithm development. In the following sections, we first propose a BCD-based algorithm for non-colluding Willies, and then modify it to address the collusion detection problem.
III Proposed BCD Algorithm